OpenStudy (anonymous):

factorise and simplify x-2/2x-4

OpenStudy (aroub):

=x-2/2(x-2) =1/2

OpenStudy (anonymous):

take 2 as common factor in the denominator then you have a cancelling step to reach to what aroub got.

OpenStudy (anonymous):

what aroub says but where x is not 2 needs to be put in as a conditioning statement so you are not dividing by zero